### v1

# An open framework for archival, reproducible, and transparent science

An open framework for archival, reproducible, and transparent science
Sabar Dasgupta Department of Bioengineering Paul Nuyujukian Department of Bioengineering Department of Neurosurgery Department of Electrical Engineering Wu Tsai Neurosciences Institute Stanford Bio-X
1 Abstract
Digital computational outputs are now ubiquitous in the research workflow and the way in which these data are stored and cataloged is becoming more standardized across fields of research. However, even with accessible data and code, the barrier to recreating figures and reproducing scientific findings remains high. What is generally missing is the computational environment and associated pipelines in which the data and code are executed to generate figures. The archival, reproducible, and transparent science (ARTS) open framework incorporates containers, version control systems, and persistent archives through which all data, code, and figures related to a research project can be stored together, easily recreated, and serve as an accessible platform for long-term sharing and validation. If the underlying principles behind this framework are broadly adopted, it will improve the reproducibility and transparency of research.
2 Introduction
Over the past few decades, researchers have recognized the large effort required to replicate experiments and the rate at which these replication studies produce negative results [28]. Dubbed a “replicability crisis” [43], the current paradigm for communicating experimental methods and distributing research outputs impedes future collaboration and the development of public trust in findings. To improve replicability, both the processes for data collection and those for creating reproducible analyses must be improved. Reproducibility is the primary focus of this manuscript with a goal of minimizing the time, effort, and expertise required to analyze data from a study beyond its published manuscript [25].
In this writing and in the context of computational workflows that perform analysis with code on data to generate results, reproducibility is defined as the measure of ease by which one can run an equivalent workflow with original data to reproduce published results. Then, replicability is the measure of ease by which an equivalent workflow can be run with independently collected data to verify published results [8]. Research that is replicable is reproducible by default so long as the original data is available and usable, but the converse is not necessarily true. Replicability sets a higher bar due to the cost associated with collecting data to repeat experiments. At a minimum, research findings should be easily reproducible to foster collaboration and knowledge access, lower the barrier to building on an existing piece of work, and comply with new funding mandates.
By the end of of 2025, research funded by the US federal government must make data accessible at the time of publishing [54, 49]. Each funding agency has their own specific policies for data sharing [37, 40, 50, 32, 13, 51]. While these policies offer guidance for researchers pursuing collaboration and open science, they do not outline how data, and more importantly, entire workflows should be shared and made accessible. The details around implementing computational best practices are left up to researchers, which is generally outside their scope of expertise and takes the focus away from core scientific work.
There are a number of standards put out by the scientific community to address these concerns ranging from generic (and not necessarily research-specific) to highly focused on one field of research. Data standards include the FAIR Guiding Principles [56] and DataFed [12]. While FAIR is a good set of principles surrounding the sharing of data, it alone does not assure reproducibility as it does not consider computation. Standards for reproducibility must advise on not only how data is stored, but also how it is manipulated and presented in analysis.
Figure 1: The research workflow involves collecting raw data, transforming it into derivative data, and producing experimental outputs using code. Ideally, all pieces of the workflow should be archived as research artifacts—the data, computation, environment, experiment outputs, and associated manuscript. A signature of each piece of data can then be kept track of using a timestamp authority to keep a trusted historical record.
Frameworks such as Spyglass (for neuroscience) or Galaxy (for bioinformatics) recommend field-specific workflows with associated software and are an excellent step towards integrating various software frameworks, but can be fragile as they rely on the permanence and compatibility of external services and APIs [30, 24]. Journals are starting to place a focus on software quality and availability and are recommending ways in which scientific code should be written in order to ensure accessibility [29]. Major journals such as Nature Scientific Data also have specific requirements and recommendations for deposition of data and code, but fall short of promoting accessible pipeline reproducibility [45]. Specifically, Nature and Science now prohibit the use of GitHub as a code deposition framework because they can be altered at a later time [35, 44] and Cell strongly recommends archives that support digital longevity [4]. New SaaS tools such as Code Ocean have been developed to address these issues and offer reproducibility and analysis pipeline visualization as a service [7]. Code Ocean uses Amazon as a cloud provider to reproduce analyses and allows exporting of the files that define the workflow environment—Containerfiles—so that they can be run elsewhere. While this is a useful tool for increasing reproducibility for the journal review process [5], it does not define an open standard and is a centralized, commercial solution that is not accessible during the research lifecycle. The calkit Python package is another useful reproducibility tool and acts as a wrapper around some of the needed technologies discussed later in this manuscript. [3]. These existing environment solutions do not yet define an open framework that aims to minimize compliance burden and the time and effort required to reproduce a research workflow many years down the line. Neither highly opinionated field-specific frameworks nor existing generic research workflow aids provide end-to-end support at the computational environment level for a wide array of fields.
The medal scale (pertaining to life sciences) [25] recommends automating dependency installation and describes a gold standard where analyses are reproducible with a signal command. Unless these dependencies are versioned and running in a consistent environment, there is no guarantee that they will produce deterministic outputs or even run without error in the future. While reproduction effort is a crucial metric, a further key consideration is how far into the future the workflow is able to be run in one step. To truly meet this gold standard the complete software pipelines associated with experiment analyses must be captured as a container image—a binary file that bundles all dependencies, including the operating system, needed to recreate the experiment outputs.
Until recently, reproducing pipelines encompassing a workflow was difficult, if not impossible, without locking in to specific computational platforms. Researchers concerned with reproducibility will often specify versioned dependencies and the operating system they used to run their code, but this may miss hidden dependencies, and further burdens researchers attempting to reproduce findings with the task of recreating the computational environment. For example, code run in a Python virtual environment installed with the same pinned dependencies may produce different outputs depending on the operating system version or the versions of other packages installed on the operating system. Further, these approaches assume that software library repositories such as PyPI for for Python or CRAN for R will continue to exist decades into the future. Containers, now “a pillar of scientific computing”, offer a solution to environment management that allows checkpointing entire workflows while being relatively operating system agnostic across local and cloud settings [38]. This manuscript describes a framework which aims to be a flexible blueprint that is compatible with as many existing systems as possible. Specific recommendations are made that allow pipeline reproducibility and define a compatibility layer for those looking to verify results and collaborate.
3 Framework Definition
The Archival, Reproducible, and Transparent Science (ARTS) open framework for accessibly sharing research involving data collection and computational analyses is defined here. The goal of this framework is to provide those performing computational analyses that produce research outputs (digital artifacts) a set of best practices for tracking the evolution of a computational research effort, archiving data and results, sharing said artifacts, and making it simple for others to reproduce results and work within a consistent analysis environment. The framework can be used collaboratively, incorporating edits and sharing across multiple individuals and groups. Pieces of the ARTS open framework may also be omitted to define best practices for partial workflows—for example, archiving just datasets, archiving code that operates on already-archived datasets, or archiving a standalone computational environment that is used in many experiments.
For a research output that includes data and analysis code to be maximally compatible with the framework, it must include the following:
  • Appropriate data and software licenses
  • Research outputs deposited in an accessible, persistent, and trusted archive
  • All analysis code tracked under version control
  • Data and code provenance documented using trusted timestamping
  • Clear documentation
  • Configuration file and corresponding file structure
  • Container definition and associated image
The points outlined above and described in detail below should ideally be incorporated early into the research lifecycle. That way, when it is time to create a deposition and archive the research output, adoption of the framework does not pose an undue burden on the submitter. Further, having a portable deposition that can be uploaded to multiple archives makes research outputs more resilient.
3.1 Archives
One goal of archiving experimental results is to make all aspects of analysis available to future experimenters for decades to come: processing data, running models, and generating figures. In order to achieve this, it is vital that the archive service housing data will likely be in existence and freely accessible as far as possible into the future, or that depositions are able to be archived in a decentralized manner. When choosing an archive, there a few key considerations to make:
An ideal archive must be publicly accessible, meaning that repositories are available for download without paywalls or logins [45]. Accessible archives must also support open licenses and allow works to be searchable by included metadata like keywords. In the ideal case, an accessible repository would enable each individual file in a deposition to be individually addressable via a static URL, with a common URL prefix for the entire deposition, and the file contents accessible under the same name and directory structure as uploaded by the researcher.
Archives must allow for persistent storage of artifacts. Once results are published, they should remain immutable and allowances for future updates or edits must result in an immutable history that includes the original publication. If history is not immutable, trusted timestamping information must be provided as described below. Persistent archives also must support persistent identifiers of some kind. Ideally this should be a digital object identifier (DOI), but a persistent URL (PURL) that is not a DOI may suffice if the service providing the DOI does not support other aspects of the framework. In these cases, associated articles should be published with a DOI and include a PURL that references the packaged experiment artifacts. Persistent archive services should also store a copy of data across multiple geographic regions so that they are resilient to external events that might result in data loss such as fires or flooding.
An ideal archive should also be trusted. Requirements for trust cannot be singularly defined in a broad sense, but at a minimum, an archive service should be established and well supported with long-term financial support and a dedicated community. Archives backed by nonprofit institutions are also preferable as they are less likely to change their policies due to economic pressures. Providing for an optional confidential review process of artifacts also promotes trust in the archive and the data it houses [45].
Figure 2: No popular research archive currently supports all aspects of the ARTS open framework, but Zenodo supports all aspects outside of hosting a container registry and allowing direct git clones of analysis code . Each column (besides storage limits) defines a question with checks and X’s indicating yes and no, respectively. Columns definitions are as follows: 3rd party upload: Does the archive allow anyone to upload data? Immutable: Once published, is deposition content locked from edits? DOI: Does the archive supply a Digital Object Identifier for uploads? PURL: Does the archive supply a persistent URL for uploads? Geographic redundancy: Is data uploaded to the archive backed up in separate geographic locations? Reviews: Does the archive provide a process for privately and confidentially sharing uploads to reviewers before publishing? Nonprofit: Does the maintainer of the archive operate under nonprofit status? Storage limits: Maximum file and deposition size as well as any other deposition limits. Bare repo: Does the archive allow depositions to be directly git cloned? Container pull: Does the archive implement a container registry to allow images to be pulled directly by container engines?
Some academic institutions offer in-house archive solutions: the Dataverse Project originally developed at Harvard, DSpace developed between MIT and HP, and the Stanford Digital Repository (SDR) are some examples of archives that promote open access and provide persistent identifiers. DSpace and Dataverse are open-source tools with instances hosted by many partner organizations while the SDR is only available to Stanford affiliates. There are also popular and journal-recommended third-party solutions such as Zenodo, Dryad, or figshare if a field-specific data repository cannot be found. Dryad allows up to 300GB of storage per submission with a publication fee, while Zenodo and figshare do not require a fee to upload data, but have stricter storage limits.
A final consideration when archiving is the ease of which content can be uploaded by researchers and then downloaded by interested parties. The upload process should be as simple and automated as possible and ideally mirror how researchers already transfer their data between storage options. The Internet Archive, for example, supports uploads of data with simple file management tools like rclone [27], while the Dataverse Project provides ingestion plugins for popular data management tools like iRODS, Globus, and DataLad [48]. In order to support accessible downloads, archives should ideally preserve artifact file structure in URL paths so that deposition files can be cloned via git as discussed in the next section. Stanford’s digital repository exposes raw file structure in this way and the current best public alternative for supporting this feature is the Internet Archive, which unfortunately does not provide immutability or DOI support. Zenodo is overall the best choice for ARTS-compatible deposition at the moment even though it obfuscates individual file paths as a zip archive. A gold standard archive service would also host a container registry in order to centralize and make accessible computational environments used to process deposited data. A full visual comparison of the discussed archive options is provided in Figure 2.
3.2 Software version control
Git is a “fast, scalable, distributed revision control system” [18] and is the most popular among version control tools which allow researchers to keep a record of experimental code and supporting files [38]. Like the Linux kernel, Git is licensed mostly under the open source GPLv2 license and is available without cost [17]. Git is great for keeping track of code and other lightweight, human readable files, but is generally not advised for use and tracking of large data files.
Experiment analysis code directories should be initialized as git repositories using the git init command and changes kept track over time as commits using git add and git commit. Researchers should be careful to not commit sensitive information such as API keys and secrets. Local environment files containing these values may be added to a .gitignore file so that git never commits them.
Git offers the git clone command to copy an entire git tree (files and history) into a new location. Accessing experiment code should be done by running this command and is often done by supplying a link to a git-hosting service such as GitHub. This for-profit service, now owned by Microsoft, does not provide persistent URLs or immutable repositories, but has made a commitment to archiving open source repos [21], has made specific archival efforts [23], and provides integrations for archiving repositories to Zenodo and figshare [22]. That said, repository owners on GitHub may delete the repository or edit its history, breaking published links and calling provenance into question.
Ideally, the code should be directly downloadable from the archive using git clone. This is possible by placing a bare git repo (the .git folder created with git init --bare) at the top-level of a desired repository to share. Most repository web servers will not support the Git HTTP smart service protocol extensions [20], and thus repositories should also be prepared with the git update-server-info command before deposition [19]. This functionality is currently supported by the Internet Archive and SDR.
3.3 Containers
The backbone of reproducibility in the ARTS open framework is the packaging of the workflow inside a container. Containers offer the promise of executing the same environment on multiple host platforms.
Containers are executed in and managed by container engines. No specific container engine is prescribed, but engines supporting the Open Container Initiative (OCI)-compatible images such as Podman or Docker are recommended. This ensures that as the technology evolves, there will be a way to run the container in the future even if the specific runtime used by the original authors is no longer supported.
Containers are specified by Containerfiles which define any relevant predefined base environment, commands to run and file structures to set up in that environment, and a jumping off point into the environment (entrypoint). Containers can take an environment variables file as an input during runtime—in this case the computational configuration config.env file defined earlier. Files in the local file system can also be made accessible within the container by using volumes—essentially mounting a directory from the local file system within the container.
To fully conform with the ARTS open framework, the provided container should support the following workflows:
  • run analysis and generate static research outputs (e.g., figure plots)
  • run interactive development environment (e.g., jupyter notebook server)
  • run interactive command line interface (e.g., ipython kernel)
Ideally, these scripts should be implemented early in the research lifecycle to provide convenience during analysis instead of acting as a barrier to submission. Additional commands provided with the container to perform actions like generating interactive plots and complete manuscript figures, or running data collection pipelines should also be included when possible.
Finally, container images defining the computation environment must be published along with other research outputs. Container images are compiled, executable packages that are the end result of building a Containerfile. It is recommended to compile container images on an x86_64 host machine and release it for that architecture at a minimum, but other host architecture images may be used and released as well (e.g., i386, arm64).
3.4 Trusted timestamping
Trusted timestamping (RFC3161/RFC5816) is a technique that provides secure tracking of creation and modification dates for any digital content including data and code [38]. Since data and code have different life cycles, the trusted timestamping process differs between them.
Data should be timestamped during the collection process. This can be achieved using the scripts available in the Trusted Timestamping Framework for Scientific Research [2] or manually with the associated web tool [1]. The framework generates a timestamps.json file which should be included at the top-level directory where timestamped data is present. An example timestamped dataset is provided for reference [39].
By applying trusted timestamping to a git code repository, the commit history can be validated even if the code hosting service does not offer immutable repositories. The process of adding trusted timestamping to experimental git repositories is simple using the framework referenced above. The relevant post-commit file is added to the .git directory of the repository and all subsequent commits result in a timestamp record being generated and stored in a publicly available record. This is a straightforward and open-source mechanism to create an electronic lab notebook.
3.5 Documentation
At the very least, basic documentation included as a README file should be present at the top level of the deposition directory detailing an overview of the dataset and its formats, instructions on how to run containerized code, and information on the outputs produced by this code. Docstrings and concise, descriptive comments within code files are also highly recommended.
From the standpoint of reproducibility, it is most important that the instructions to setup and run analysis on the provided data are clear and simple. Recommended container dependencies (e.g., docker/podman version, tested architectures, etc.) should be provided as well as any scripts for convenience.
From the standpoint of replicability, clear methodology defining the data collection process should also be present. The fine details of this process, if included in research articles, are often left to supplementary sections and not usually accompanied by data collection pipelines. Artifact documentation should include technical details around data collection such as experimental setup, materials (including specific hardware) needed, and any related code.
3.6 Configuration and file structure
An environment variables file named config.env must be present at the top level folder of the deposition or under a folder named arts/ to be compatible with the ARTS open framework. This file defines where code, data, and research outputs are located, locations for other required files, and any other parameters that may be passed to analysis or data collection code. All other files and folders, while recommended to include when relevant, are optional and may be omitted.
The following is an example default environment variables file specifying a recommended set of file and folder locations for an ARTS-compatible research output. Path variables use Unix-like syntax and should specify relative paths with respect to the top-level directory:
ARTS_RAW_DATA_PATH=data/raw/
ARTS_DERIV_DATA_PATH=data/deriv/
ARTS_CODE_PATH=code/
ARTS_ENV_PATH=env/
ARTS_OUTPUT_PATH=output/
ARTS_LICENSE_PATH=LICENSE
ARTS_README_PATH=README.md
ARTS_RUN_CMD=./run.sh
ARTS_SETUP_CMD=./setup.sh
Researchers are responsible for using these environment variables in their code so that changing them has an affect on analysis reruns. Omitting an environment variable or setting it to an empty string (e.g., ARTS_CODE_PATH= ) indicates that corresponding file or folder is not present in the deposition. In some circumstances, researchers may also want to provide a PURL or other resolvable value in path variable. For example, if new features are being derived from an existing raw dataset, the ARTS_RAW_DATA_PATH may be set to an external DOI or PURL where the data is accessible, so long as the setup and run commands support using an external URL.
In the case where a study includes multiple git repositories, each git repository can be included in a separate folder under the code directory. Then, if a top-level git repository is defined for the deposition, the code/ directory can be placed in a .gitignore file to avoid clashing git repositories. Alternatively, git submodules may be used.
Figure 3: Minimal example file structure adhering to the ARTS open framework.
The presence of a config.env greatly simplifies the reproduction process by allowing validators to rerun analyses and recreate the output/ directory with the one-line ARTS_RUN_CMD. Replication is also largely simplified for results that are reproducible through the ARTS open framework. After running a new experiment to collect independent data, the value for ARTS_RAW_DATA_PATH can be updated to run analyses on these data and complete the replication. Incorporating top-level environment variables during the analysis and data collection processes is also a recommended best practice as it provides a unified interface for parameter tuning and makes code more accessible.
3.7 Licensing
Licenses define what rights you give others to use, modify, and share your code and data. By default in most jurisdictions, all rights to use novel creations are reserved and so it is necessary to spell out what uses are allowed [47]. Omitting an appropriate license from an otherwise accessible deposition will make it unusable. Open source code and open content licenses, as opposed to proprietary or noncommercial licenses, give free access to use, modify, and share content [42, 15]. Common choices for open source code licenses include permissive licenses such as BSD-3, MIT, and Apache-2.0. Share-alike or “copyleft” licenses such as the GPL family are also a popular choice as they require modified code to be distributed with a similar license and therefore encourage return contributions from collaborators [14].
As for data and other content such as documentation, Creative Commons (CC) provides three licenses that are widely used and recommended. Open Data Commons (ODC) also provides a set of similar open licenses that are aimed specifically at data, but may not be suitable for other content [41]. The most recent versions of the CC licenses are recommended over ODC licenses as they provide broader coverage and permissions [9]. For placing data in the public domain without restriction, the CC0 license should be used, and is required for work directly generated by US federal agencies and staff. If attribution to the original authors is required during redistribution, the CC BY license is suitable. And if an attribution and share-alike clause is desired, there is the CC BY-SA license [10]. Unfortunately, none of these licenses provide terms for data provenance beyond attribution. For example, even a dataset released under the CC BY-SA license can be downloaded, slightly modified, and reuploaded as a purported exact copy of the original so long as it is released with attribution information and a copy of the license. Therefore, it is important to release attestable provenance information along with datasets as described in the trusted timestamping section.
Figure 4: Open licenses apply to software and data and can require attribution, or openness of derivative work. Permissive licenses assign copyright, but give users broad usage and distribution rights whereas works placed in the public domain release copyright from the original author. Creative common licenses are popular for data and content, but open-source code is more often licensed under one of the other listed licenses.
When choosing a license, there are a few important considerations to make [14]. There may be requirements set by government or foundational funding sources requiring research outputs be released under specific terms [53, 16, 26]. Specific archives may also have license requirements—for example, Dryad places all depositions under a CC0 license. After fulfilling any obligations, the intended audience of the code and data should be considered. Results that are intended to be built upon in a research setting may be best served under a share-alike license to encourage future openness. To maximize reach across commercial settings, placing code under permissive licenses and data in the public domain is recommended. Conversely, a noncommercial license may be used if the goal is to discourage commercial use. Dual licensing may also be an option for authors that wish to retain the ability to license their work commercially under stricter terms [52]. For research outputs, it is generally best to use a license that requires attribution at a minimum to preserve the historical record. An overview of recommended open licenses is provided in Figure 4.
4 Recommendations
The following recommendations apply to specific groups of stakeholders involved with running, archiving, sharing, and funding research.
4.1 Researchers
Researchers should fully document their workflows if they do not do so already. This includes spelling out all versions used, including packages and libraries used in code, programming languages, and operating systems tested on. Shell scripts that define and run the workflow should also be provided. Finally, providing a container image that bundles all aspects of the experiment is necessary in order to create a community around reproducing results and sharing scientific knowledge as broadly as possible. Implementing the ARTS open framework is a good way to achieve this and for well-designed computational studies, should primarily require specifying a config.env file and modifying associated top-level scripts.
4.2 Archivists
Those who administer deposition archives should make sure that they have at least the minimal feature set described by Figure 2. Except for hosting container registry endpoints, a few archives are close to being full-featured in this regard. Zenodo and the Dataverse Project should expose file structure in URLs to support git repo cloning. The Internet Archive could be a suitable archive for research projects if it added support for immutable releases and temporary private uploads for the purpose of peer review. Stanford’s Digital Repository would also support all aspects of the standard if it opened up submissions to external collaborators. Building in useful provenance tools like trusted timestamping also moves some of the work off of researchers and increases the reputability of the archive. All major archiving platforms should support pulls for OCI containers.
4.3 Publishers
Publishers and those involved with the scientific review process should also require that the research they are distributing conforms to reproducibility best practices. The peer review process for submissions that include code should involve having an independent recreation of relevant figures from referenced data. A number of Nature journals have taken a step in this direction by integrating and encouraging the use of Code Ocean for reviews [34, 33]. Publishers, especially those already partnered with archive services, should also be aware of the opportunity to implement automated ”clearing house” services that run the deposited container to validate figure generation from data for ARTS-compatible submissions. Encouraging third-party figure recreation in this manner sets a high bar for computational reproducibility.
4.4 Funding agencies and policy makers
Many US funding agencies are taking steps to ensure that data generated from sponsored awards be made public at the time of publication [36]. While this is an excellent first step, simply having the data released does not mean the data is usable or accessible. Code and the computational environment needed to generate the outputs of research could also be required to be made public in the same manner, and this would significantly increase the accessibility of research works. The ARTS framework would be an example framework compliant with both the letter and the spirit of such requirements.
5 Example Implementation: Watch Calibration
As a simple example experiment and computational analysis to demonstrate how the ARTS open framework is implemented, audio from a Chinese standard movement [6, 55] watch was recorded with a laptop microphone in order to determine the clock drift and calibrate the watch. This experiment was chosen since it highlights all aspects of the framework while offering an accessible way to replicate the experiment by generating a new dataset to anyone with a computer microphone and mechanical watch.
A 2813 automatic mechanical movement was recorded with a laptop by placing the movement and case on the interior of the laptop above the headphone jack where the microphone is located. Other watch movements and recording equipment may be used, but when recording with a laptop it’s important to place the watch directly on top of the laptop microphone and record in a quiet environment. Audio data was recorded at 44100 Hz using the sounddevice Python package. The experiment code also offers the ability to generate a dataset using the librosa.clicks method [31]. Derivative data in the form of audio windows containing each clock tick and their respective peak onset time were output to the ARTS_DERIV_DATA_PATH folder.
Figure 5: Comparison of drift values across each dataset recorded (W241130; 250227) and generated (250409; 250310_noisy) for all possible analysis parameters. Drifts are fairly consistent for recordings on the left with high SNR. Data with more noise maybe need filtering for a reasonable drift calculation (250410_noisy) or filtering may result in additional noise (250227).
To calculate the drift of the watch movement over time, the audio signal peaks were picked out using scipy.signal.find_peaks. Clock tick windows were then defined and adjacent windows were correlated with one another to adjust for noise in calculating the local maxima. Prominence parameters for find_peaks were tuned to work using an audio signal with a high signal-to-noise ration, so noisy audio data may require additional filtering or optional usage of an envelope calculation included as parameters in the analysis.
Figure 6: Visualization of difference between a tick at the end of the recording (dashed line) and a corresponding tick from an ideal generated source (dotted line) as in the case of no filtering. The difference between these tick times corresponds to the first row in Figure 5
The container containing the analysis can be run via Podman or Docker using the ./run.sh script available in the deposition [11]. The run script provides a command to regenerate figures (./run.sh generate-figures), launch a jupyter notebook (./run.sh jupyter), or launch an ipython interactive shell (./run.sh ipython). The recording process can also be rerun using the collect_data.sh script to collect new data using a system microphone or generate new data using librosa. The container image may also be regenerated with ./run.sh save-image.
6 Limitations and Future Work
6.1 Limitations
The ARTS open framework is not a silver bullet for research reproducibility and replicability, but simply a set of flexible best practices. It is not a replacement for careful data collection, statistical rigor, or clear explanation of methods. The framework is dependent on a number of other technologies, notably: git (although it may be extensible to other version control systems), the OCI’s container specifications, and chosen archive services. It also does not yet adequately address workflows that work with large raw data on the order of terabytes or more per experiment or those that involve specific or proprietary hardware.
And although the ARTS open framework offers wide support across compute environments, steps to perform GPU workflows or hybrid workflows that span local and cloud compute clusters are not prescribed in this manuscript or accompanying code. Further, container images generated on x86 host systems will only reliably run on other x86 systems and container support across various computing architectures may be limited [46].
It should also be noted that container images for scientific computing tend to be quite large (on the order of 1-10 GB depending on the included dependencies). Archives serving these files can take advantage of layer caching—the process of sharing reusable blocks of containers that share dependencies—to greatly reduce the total storage needed to host container images.
6.2 Future Work
ARTS is a generic framework, but to make life easy for researchers in their respective fields, it’s important to have opinionated frameworks. One signal processing example experiment using the framework is provided at the time of writing, but for the ARTS open framework to be most useful, there must be further examples created adhering to the framework that detail specific data collection pipelines and shared environments for analyses using programming languages besides Python, such as R and MATLAB. Clear examples on how to archive large data on the order of terabytes with trusted timestamping and version control need also be added.
To improve adoption of archiving computation environments among publishers, an automated figure validation framework is needed which will run the ARTS_RUN_CMD using the provided container image. This service should be exposed initially as an open source infrastructure as code (IaC) template to encourage standardization across publishers.
7 Conclusion
Science, especially that which is funded by the public, is an asset that should be accessible to the public including other researchers. Although new standards and guidelines make the outputs of scientific endeavors available, they are often not easily usable, even to other experts in the field. This large burden of making results reproducible currently falls on researchers and highlights the importance of tools that are compatible with already-used workflows.
To bridge this knowledge gap requires education on both sides. Researchers and interested end users need to learn at least the high-level DevOps principles necessary to output their work in an accessible format and run those outputs. Fluency with these techniques are often left out of university curricula and treated as professional skills, but they are now core to modern science.
Given that no current research artifact archive meets all recommendations of the ARTS open framework, archive services are encouraged to preserve deposition file structure in URLs and move towards supporting container registries that house the container images used to validate ARTS-compatible research outputs. Implementing these features will allow for automatic recreation of figures and validation of computational results.
By providing a framework for reproducible workflows, the hope is to enable other researchers to further increase the replicability of their work and encourage the focus of replication to be on sound scientific methods for robust data collection and not interpretation of analysis.